OpenStudy (anonymous):

how do you solve 4*n-3=27? i have to find n

OpenStudy (anonymous):

n= 5

OpenStudy (anonymous):

First you add 3 to both sides and then divde by 4 and the answer is 11

OpenStudy (anonymous):

Because 4*5 = 20 20 - 3 =27

OpenStudy (anonymous):

are you sure you know what you are doing?

OpenStudy (anonymous):

and you can add 3 to 27 = 20 then divide 20/4 = 5

OpenStudy (anonymous):

I'm sorry but i did the problem right not you

OpenStudy (freckles):

do you know 20-3=17 actually?

OpenStudy (anonymous):

whos right? im very confused

OpenStudy (anonymous):

Ya made a fool of my self.

OpenStudy (freckles):

try adding 3 on both sides as a first step then divide both sides by 4

OpenStudy (anonymous):

I must be right

OpenStudy (anonymous):

n = 7.5

OpenStudy (anonymous):

7.5*4 = 30

OpenStudy (anonymous):

30-3 = 27

OpenStudy (anonymous):

add 3 to the 4 or 27?

OpenStudy (anonymous):

Don't listen to kevin

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

wait how did you get 11?

OpenStudy (anonymous):

i got what kevin got i must have done it wrong

OpenStudy (anonymous):

ok thanks anyways

OpenStudy (anonymous):

Your welcome